¡ Obtaining Files By Anonymous FTP ¿ (Adapted from RFC 1175)

Files may be obtained with the FTP program in conjunction with an ANON-YMOUS login. Versions of the FTP program may vary from system to system, so the commands shown in this example may need to be modified to work on your system.

% ftp nic.funet.fi ¡== Use the FTP program to

connect to nic.funet.fi

Connected to nic.funet.fi
220 nic.funet.fi FTP server (Version 4.471 Feb 21 19:27:02 EET) ready.

The system should respond with a message to indicate that a connection has been made. Users on a Unix system will probably be prompted for a login name. Type in ”anonymous” as in the example below:

Name (nic.funet.fi:user): anonymous
331 Guest login ok, send ident (your email address) as password.

Password: ¡== Type in ¡your-email-address¿ at the password prompt Other systems may require the use of a ”login” or ”user” command to send

the username to the server computer. Users unsure of the command should contact a local site representative for the specific commands.

After the username and password are sent to the system, a message to indicate that the login has been made successfully should appear:

230-Guest login ok, access restrictions apply.
230-Local time is now Sat May 11 11:42:12 1991
230-There are constant overload problems on our line to USA...
230-  Be careful with "cd ..".  See  /README for info about how
230-  symlinks might affect your view of, and movement in archive.
230-We have special access features, see file /README
230 It was last updated Thu Feb 21 19:30:35 1991 - 78.6 days ago

The user then connects to the directory in which the document to be retrieved resides. This is done with the cd command:

ftp> cd pub/misc/computer-english

250 CWD command successful.

The user should now be connected to the computer-english directory. The ”dir” or ”ls” command will list the files available in this directory.

ftp> dir
200 PORT command successful.

150 Opening ASCII mode data connection for .

*** At this point a list of the files in the directory

should appear **

226 Transfer completed.

The ”get” command will get any file in the directory.

ftp> get 1-Introduction.tex

200 PORT command successful.
150 Opening ASCII mode data connection for 1-Introduction.tex (4482 bytes).
226 Transfer complete.
4482 bytes received in 2 seconds (2.2 Kbytes/s)

The ”quit” command leaves the FTP program.

ftp> quit
221 Goodbye.